OpenStudy (anonymous):

plz help me with an 8th grade question

OpenStudy (anonymous):

Colton solved an equation incorrectly, as shown below: Step 1: x − 13 = 26 Step 2: x = 26 − 13 Step 3: x = 39 Which statement best explains why Step 2 is incorrect in Colton’s solution? He did not subtract 13 from 26. He did not multiply 26 by 13. He did not add 13 to 26. He did not divide 26 by 13.

OpenStudy (anonymous):

@icecreamman @iLorita @animalfan1

OpenStudy (anonymous):

plz anyone

OpenStudy (anonymous):

hmm im thinking

OpenStudy (anonymous):

thank u

OpenStudy (anonymous):

im thinking it is D. but im also thinking it is C. also...

OpenStudy (anonymous):

ok...

OpenStudy (anonymous):

well you know its not A... but to get to 39 you would have to add I think C

OpenStudy (anonymous):

yeah that might be it...

OpenStudy (anonymous):

ok so c

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

it would probablybe the right one so yeah

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

u welcome :D